### v1

# An Investigation of Robustness of LLMs in Mathematical Reasoning: Benchmarking with Mathematically-Equivalent Transformation of Advanced Mathematical Problems

Abstract
In this paper, we introduce a systematic framework beyond conventional method to assess LLMs’ mathematical‑reasoning robustness by stress‑testing them on advanced math problems that are mathematically equivalent but with linguistic and parametric variation. These transformations allow us to measure the sensitivity of LLMs to non-mathematical perturbations, thereby enabling a more accurate evaluation of their mathematical reasoning capabilities. Using this new evaluation methodology, we created PutnamGAP, a new benchmark dataset with multiple mathematically-equivalent variations of competition-level math problems. With the new dataset, we evaluate multiple families of representative LLMs and examine their robustness. Across 18 commercial and open-source models we observe sharp performance degradation on the variants. OpenAI’s flagship reasoning model, O3, scores 49 % on the originals but drops by 4 percentage points on surface variants, and by 10.5 percentage points on core-step-based variants, while smaller models fare far worse. Overall, the results show that the proposed new evaluation methodology is effective for deepening our understanding of the robustness of LLMs and generating new insights for further improving their mathematical reasoning capabilities.
Introduction
Figure 1: PutnamGAP Variants Performance Relative to the Original Set
Motivation
Modern AI systems are increasingly entrusted with tasks that hinge on reasoning rather than pattern matching. Reliable progress therefore depends on precisely measuring an LLM’s reasoning capacity and its ability to generalize beyond memorized textual surface forms. Existing math-reasoning benchmarks, however, exhibit two critical weaknesses: (i) leakage-induced score inflation, since benchmark items rapidly seep into pre-training corpora, and (ii) limited robustness coverage, because today’s datasets are too small or lack controlled transformations that probe true generalization. Addressing these weaknesses is urgent if we aim to benchmark reasoning with the same rigor demanded in safety-critical domains such as healthcare or cybersecurity.
Benchmark inflation through training leakage.
Recent studies show that public datasets, including GSM8K(Cobbe et al. 2021) and MATH (Hendrycks et al. 2021), have leaked into the web‑scale corpora used to pre‑train large language models (LLMs), artificially inflating test‑time accuracy. A leaderboard score therefore no longer guarantees genuine reasoning ability; it may merely reflect memorization of benchmark items or their solutions. Simply releasing yet another dataset postpones the problem: once its items enter future training corpora, scores climb without real progress. What is needed is a systematic method that (i) measures a model’s capacity to generalize beyond verbatim memory and (ii) can generate an unbounded supply of evaluation items, limiting future leakage.
Competition mathematics reveals the next robustness bottleneck.
Large language models (LLMs) now surpass 90% accuracy on widely‑used benchmarks such as GSM8K and MATH, prompting claims of “near‑human” numerical reasoning yet still falter on Olympiad‑style or Putnam‑level problems that intertwine multiple domains. Existing Putnam‑derived datasets are too small to expose this gap: Putnam‑AXIOM (236 originals + 52 variations) (Huang et al. 2025), and PutnamBench (640 formalized theorems) (Tsoukalas et al. 2024) remain in the hundreds, and none delivers systematic generalization and perturbations. These facts expose Weakness (i) insufficient scale and Weakness (ii) lack of controlled, systematic transformations in existing evaluations.
Generalization–and–Perturbation (GAP).
We address both leakage and robustness with a simple idea: stress‑test the model on mathematically equivalent versions of the same problem. For a problem $x$ with solution set $S(x)$ and an LLM $f$, robustness is the expected accuracy when $x$ is transformed by a family $\mathcal{T}$ of equivalence‑preserving operators (§ 2 gives the formal definition). We partition $\mathcal{T}$ into $\mathcal{T}_{\text{surf}}$ (surface renames that alter symbol salience) and $\mathcal{T}_{\text{para}}$ (kernel rewrites that preserve the same proof steps while changing the scenario and parameters). This GAP framework (i) creates an infinite stream of unseen test items, mitigating future contamination, and (ii) quantifies how far a model can generalize beyond memorized surface forms.
PutnamGAP: instantiating GAP on 85 years of problems.
We instantiate GAP on every William Lowell Putnam Competition problem from 1938–2024 (1 051 originals) and expand each item into five variants—four surface renames and one kernel rewrite—obtaining 6 306 stress‑test questions.A two‑stage QA pass—15 rounds of o3 self‑review plus a 10% spot‑check found no substantive errors.
Headline results.
Across ⟨17⟩ models, as shown in 8, all of them suffer from both simple renaming and step-based rewrites. OpenAI’s o3 scores ⟨49⟩% on original statements but loses ⟨4⟩ pp under surface renames and ⟨10.5⟩ pp under kernel rewrites. These drops confirm that high leaderboard scores can collapse when cosmetic or structural perturbations are applied—precisely the effect that data leakage masks.
Contributions.
(1) We introduce GAP, a general framework for measuring robustness via mathematically equivalent transformations. (2) We release PutnamGAP, the first 6k‑scale competition benchmark that systematically disentangles surface‑level and structural generalization while limiting future leakage. (3) We provide the first comprehensive robustness baseline across seventeen LLMs, plus an open‑source evaluation stack.
Method
We start from a curated set of $N$ canonical items
$\mathcal{P}=\bigl{\{}(x_{i},y_{i},\pi_{i})\bigr{\}}_{i=1}^{N},$
where $x_{i}$ is a problem statement, $y_{i}$ is its reference answer(s), and $\pi_{i}$ an unreleased expert solution path used internally for safe variant generation.
Model interface.
A language model $f_{\theta}$ receives a prompt $x$ and returns $\hat{y}=f_{\theta}(x)$, which an automatic checker maps to a binary label $z=\operatorname{grade}(\hat{y},y)\in\{0,1\}$.
Variant families.
For every $x_{i}$ we later apply two disjoint transformation super‑families (defined in the next section but left unchanged here):
$\mathcal{T}_{i}^{\textsf{surf}}$ $\;(\text{$K_{surf}$ surface variants}),$ (1)
$\mathcal{T}_{i}^{\textsf{para}}$ $\;(\text{$K_{core}$ parametric variant}).$ (2)
Each surface transformation $\tau$ returns a new statement $x_{i}^{(\tau)}=\tau(x_{i})$ that preserves semantic correctness of $y_{i}$. For parametric variations, $y_{i}$ is transformed as well to match $\tau(x_{i})$
Evaluation matrix.
The Cartesian product $\mathcal{D}=\{(i,\tau)\mid i\leq N,\;\tau\in\mathcal{T}_{i}^{\textsf{surf}}\cup\mathcal{T}_{i}^{\textsf{para}}\cup\{\mathrm{id}\}\}$ contains $N\times(K+1)$ aligned items (original + $K$ variants per source, $K=K_{surf}+K_{para}$). Running $f_{\theta}$ on every pair populates a binary matrix $\mathbf{Z}\in\{0,1\}^{N\times(K+1)}.$ From the first column we extract the easy vector $\mathbf{e}(\theta)\in\{0,1\}^{N},$ while the remaining columns feed family‑specific aggregates:
$\mathbf{h}^{\textsf{surf}}(\theta)=\text{maj}\bigl{(}\mathbf{Z}_{[:,\,\textsf{surf}]}\bigr{)},\quad\mathbf{h}^{\textsf{para}}(\theta)=\mathbf{Z}_{[:,\,\textsf{para}]}\!.$
The set of surface variants can be changed based on specific tasks.
Robustness Metric
Let $\mathbf{e},\mathbf{h}\!\in\!\{0,1\}^{N}$ be per-item correctness on the easy and hard sets. In this work, easy sets refer to the original and hard sets refer to the variants. Let $p_{e},p_{h}$ be their Jeffreys-smoothed accuracies, and define the pooled standard deviation $\sigma=\sqrt{\tfrac{1}{2}\bigl{[}p_{e}(1-p_{e})+p_{h}(1-p_{h})\bigr{]}}$. We measure robustness with
$R(\mathbf{e},\mathbf{h})=\frac{1}{N}\sum_{j=1}^{N}\exp\!\Bigl{(}-\tfrac{e_{j}-h_{j}}{\sigma}\Bigr{)}.$
The exponent converts accuracy drops (in $\sigma$ units) into a multiplicative penalty in $(0,1]$: $R=1$ means perfect invariance, while smaller values indicate greater brittleness. Full derivation, statistical justification, and design discussion are in Appendix B.
Robustness scores.
Using the robustness metric, we report
$R_{\textsf{surf}}$ $=R\bigl{(}\mathbf{e},\mathbf{h}^{\textsf{surf}}\bigr{)},\quad R_{\textsf{core}}=R\bigl{(}\mathbf{e},\mathbf{h}^{\textsf{para}}\bigr{)},$
$R_{\textsf{global}}$ $=\sqrt{R_{\textsf{surf}}\,R_{\textsf{para}}}.$ (3)
Only these vectors are needed—the content of the Transformation Families subsection itself is left intact.
Reproducibility.
An open‑sourced script will be released after review.
Transformation Families
The proposed general robustness measures can work for any variations. As a first step in exploring this new evaluation methodology, we propose and study five aligned variants— four surface renamings that perturb only symbol names, and one core‑step instance that perturbs numeric slots while preserving the reasoning chain. This section details the synthesis pipelines;
Surface renaming family
We want to know whether a model recognizes an argument because it has truly abstracted the pattern or merely because it memorizes suggestive identifier strings. Therefore we systematically replace each token tagged var or param; all constants of category sci_const remain untouched.
Automated pipeline
  1. 1.
    Proposal. A single call to o3 receives the token role (“free variable” or “fixed parameter”) and the surrounding textual context, and returns a candidate replacement.
  2. 2.
    Collision check. A deterministic post‑validator rejects names colliding with any pre‑existing identifier in the problem.
  3. 3.
    Family tagging. The accepted string is labelled as belonging to one of four families described below.
The Five families
DL
Single descriptive phrase (populationDensity).
DLC
2–5 unrelated nouns (walnutVioletTerrace).
DLM
Misleading mathematical term (primeOrder).
GS
4–16‑char hash (xcQ7h2ZfRw9v).
Each source item thus yields $4$ surface variants; accuracy deltas per family appear in Section Results & Analysis.
Parametric variant
Symbol renaming probes only the lexical axis. To probe structural transfer, we resample numerical constants yet force the solution to reuse the original high‑level moves. In this work, we call it Kernel_Variant (KV).
Four‑stage pipeline
We convert each item into semantically-equivalent variants through a four-stage pipeline (pseudocode (Alg. 1)):
  1. 1.
    slot discovery,
  2. 2.
    template back-synthesis,
  3. 3.
    semantic-preserving renaming, and
  4. 4.
    dual-verifier screening (two-in-a-row rule).
The pipeline generates a bounded number of validated variants for each problem within a few hours on commodity hardware using the OpenAI o3 API. See Appendix C for empirical bounds and implementation details.
Algorithm 1 Repair‑and‑verify loop (excerpt)
1:input: draft variant $v_{0}$
2:for $t=1$ to $T$ do
3:  Run $J$ o3 judges $\to$ verdict vector $\mathbf{z}_{t}$
[…]
5:   accept $v_{t}$ {two‑in‑a‑row passed}
6:   break
7:  else if $\mathbf{z}_{t}=\mathbf{1}$ then
8:   keep $v_{t}$ for next round
9:  else
10:   apply LLM‑suggested patch $\to v_{t}$
11:  end if
12:end for
13: human audit 10 % of accepted variants
Implementation Overview
Code release.
To facilitate double‑blind reviewing we publish only the subset of data (100 randomly chosen examples). An automated evaluator, putnam-cli.py, receives the names of target solver model and grader model and variant type to test. Supported back‑ends are (i) any HuggingFace‐compatible checkpoint via transformers, (ii) a local vllm server, or (iii) API clients including OpenAI, Gemini, Anthropic and OpenRouter Full data and generation scripts will be released post‑decision.
Surface generation.
Renaming variants are produced on a CPU‑only node by streaming o3 API calls. A five‑stage exponential‑back‑off retry (max 5 attempts, doubling timeout each time) masks transient API latency. Processing all 1 0511\,051 items in parallel takes $\sim$15 min wall‑clock.
Core‑step generation.
Kernel variant synthesis is more expensive because of multi‑turn chain‑of‑thought reasoning: end‑to‑end runtime is $\leq\!3\,\text{h}$ for the full corpus on a single 8‑core CPU, dominated by the 15‑iteration repair‑and‑verify loop.
PutnamGAP Dataset
Data Sources, Extraction & Annotation
Our benchmark comprises all Putnam Problems 1938–2024 (N=1 051N=1\,051 items after deduplication). The front-matter of every book contains the same fair-use clause, excerpted verbatim below:
“Individual readers … are permitted to make fair use of the material, such as to copy select pages for use in teaching or research.”
This clause grants us the legal right to reproduce problems and solutions for non-commercial academic evaluation. In line with AMS policy, we distribute only machine-readable IDs and LaTeX texts; raw PDF scans remain under the original AMS license, and any further redistribution must be cleared through the Copyright Clearance Center.
Original scans are processed via a 3-stage OCR routine: (i) Manual segmentation for every question-answer pair. (ii) MathPix for formula‐aware PDF-to-LaTeX conversion followed by (iii) custom post-filters that merge multi-line expressions and fix 4.2 % residual symbol errors. Each item is manually spot-checked ($\leq$2 min per problem) to ensure semantic fidelity before variant generation. Complete corpus list, OCR accuracy study, and cleaning scripts appear in Appendix D.
Dataset Statistics
Overall scale and balance.
Figure 2: Problem Topics and Classes
The benchmark comprises 1 051 original Putnam problems from 1938–2024 and five mathematically equivalent transformations, yielding 6 306 items. Part distribution is balanced (527 A vs. 524 B), and the canonical identifier $\langle\textit{year},\textit{part}\{A,B\},\textit{index}\rangle$ provides a difficulty proxy. Using indices 121\!\!-\!2 as Easy, 343\!\!-\!4 as Medium, and 565\!\!-\!6 as Hard, the corpus contains 32.3 % Easy, 32.3 % Medium, 32.2 % Hard, plus a 3.0 % extra–hard tail (indices 7–8).
Topic coverage.
Automatic tags in _meta.tag indicate broad mathematical coverage—Algebra (641), Analysis (521), Number Theory (392), Combinatorics (286), and Geometry (239).
Statement length.
After stripping  markup, an average statement is 35.7 tokens (min 0, max 152); all transformation families alter length by <2<\!2%, confirming that robustness drops cannot be explained by verbosity alone.
Quality control.
Every item has undergone single‐pass manual validation; inter‐rater Cohen $\kappa$ is not yet available and will be released in a future update.
Experimental Setup
Model Pool & Prompting
We evaluated 18 models (see 1 or Appendix A for a complete list).All models are queried under a unified zero‑shot template. A system instruction designates the model as “an expert mathematician” and asks it to show all work, while the user message embeds the problem. See Appendix for our full prompt. We fix temperature=0, top_p=1, and max_tokens=32000 or maximum token amount available in case some models have max_tokens maximum smaller than 32000. for every run except OpenAI O-series which require temperature=1. Solutions are then re‑submitted to a second template that grades the answer: a strict proof rubric for proof items and a lenient numeric rubric for calculation items. Both grader prompts require structured JSON output containing a binary grade field plus detailed feedback. Complete prompt code and hyper‑parameter settings are available in Appendix B and the anonymous repository (loader/prompt.py).
Scoring & Auto-Grader
We partition tasks into computational and proof-based categories and evaluate them with distinct graders.
Computational. Each candidate answer is normalized (whitespace, units, LaTeX macros) and passed to two scoring paths: (i) a strict string match against the reference solution; (ii) a latent grader—an LLM prompted to return ‘‘CORRECT’’ or ‘‘INCORRECT’’ given the reference answer and a rubric that disallows partial credit. We adopt path (ii) to mitigate formatting artifacts; if the two paths disagree we mark the item for manual audit (¡1 % of cases).
Proof-based. We provide the grader with an aligned, step-by-step reference proof and ask it to assign a binary grade plus a natural-language justification. Any skipped logical step or missing citation triggers a fail. A random 10 % sample is double-checked by independent volunteers; grader precision/recall is >>97 %.
All grader prompts, rubrics, and decision thresholds are released as supplementary material after the review period.
Compute & Reproducibility
All inference were performed through publicly available APIs. Each model was queried exactly once per item with the hyper-parameters in Table 2. Runs were executed from a single Ubuntu 22.04 host (11th Gen Intel(R) Core(TM) i7-11800H @ 2.30GHz); no local GPU was used. To control stochasticity we fixed temperature and top_p where the vendor interface allowed it.
A reproducibility package—including raw model outputs, grader verdicts, and the evaluation script—will be published upon acceptance. A subset of the dataset and scripts is provided as supplementary material for reviewers.
Results & Analysis
Table 1: Model Accuracy Rates across Categories (Percent Scale)
Model Original Descriptive ($\Delta$) Confusing ($\Delta$) Misleading ($\Delta$) Garbled String ($\Delta$) Kernel Variant ($\Delta$)
claude-opus-4 26.5∗∗∗ 23.0∗∗ (–3.5) 22.2∗∗ (–4.3) 21.7∗∗ (–4.8) 21.4∗∗ (–5.1) 13.8∗∗∗ (–12.7)
claude-sonnet-4 23.0∗∗∗ 20.6∗∗∗ (–2.5) 19.8∗∗ (–3.2) 18.6∗∗ (–4.4) 18.1∗∗∗ (–4.9) 11.1∗∗∗ (–11.9)
deepseek-prover 15.5∗∗∗ 15.2∗∗∗ (–0.3) 14.0∗∗∗ (–1.5) 12.8∗∗ (–2.7) 13.7∗∗∗ (–1.8) 9.2∗∗∗ (–6.3)
gemini-2.5-flash-lite 19.8∗∗∗ 18.8∗∗∗ (–0.9) 16.1∗∗ (–3.7) 15.8∗∗ (–4.0) 15.1∗∗∗ (–4.7) 6.6∗∗∗ (–13.2)
gemini-2.5-pro 78.4∗∗∗ 75.2∗∗ (–3.1) 74.3∗∗ (–4.1) 72.8∗∗∗ (–5.6) 72.9∗∗∗ (–5.4) 63.5∗∗∗ (–14.9)
gemini-2.5-flash 42.8∗∗∗ 42.6∗∗∗ (–0.2) 39.0∗∗ (–3.8) 40.9∗∗∗ (–1.9) 37.6∗∗ (–5.2) 27.6∗∗∗ (–15.2)
gpt-4.1 24.9∗∗∗ 23.4∗∗∗ (–1.5) 21.2∗∗ (–3.7) 22.0∗∗∗ (–2.9) 22.6∗∗∗ (–2.3) 14.8∗∗∗ (–10.1)
gpt-4.1-mini 28.6∗∗∗ 26.9∗∗∗ (–1.7) 27.1∗∗∗ (–1.5) 24.0∗∗ (–4.6) 25.1∗∗ (–3.4) 19.2∗∗∗ (–9.4)
gpt-4.1-nano 8.8∗∗∗ 8.9∗∗∗ (+0.1) 6.4∗∗ (–2.4) 7.3∗∗∗ (–1.5) 6.6∗∗∗ (–2.2) 6.8∗∗∗ (–2.0)
gpt-4o 6.5∗∗∗ 6.3∗∗∗ (–0.1) 5.3∗∗∗ (–1.1) 4.7∗∗ (–1.8) 4.7∗∗ (–1.8) 3.0∗∗∗ (–3.4)
gpt-4o-mini 4.3∗∗∗ 3.5∗∗∗ (–0.8) 2.5∗∗ (–1.8) 3.3∗∗∗ (–1.0) 3.2∗∗∗ (–1.1) 1.7∗∗∗ (–2.6)
grok4 60.1∗∗∗ 59.0∗∗∗ (–1.1) 56.6∗∗∗ (–3.4) 55.9∗∗ (–4.2) 55.2∗∗ (–4.8) 45.5∗∗∗ (–14.6)
kimi-k2 27.2∗∗∗ 25.8∗∗∗ (–1.4) 23.7∗∗ (–3.5) 23.8∗∗ (–3.4) 23.4∗∗ (–3.8) 12.8∗∗∗ (–14.4)
llama-4 15.7∗∗∗ 14.5∗∗∗ (–1.2) 13.0∗∗ (–2.7) 13.1∗∗ (–2.6) 13.8∗∗∗ (–1.9) 7.3∗∗∗ (–8.4)
mistral 5.5∗∗∗ 5.5∗∗∗ (–0.1) 5.7∗∗∗ (+0.1) 4.9∗∗∗ (–0.7) 4.2∗∗∗ (–1.3) 3.9∗∗ (–1.6)
o3 51.5∗∗∗ 52.8∗∗∗ (+1.3) 47.6∗∗ (–3.8) 49.5∗∗∗ (–2.0) 46.8∗∗ (–4.7) 38.6∗∗∗ (–12.9)
o4-mini 41.5∗∗∗ 43.0∗∗∗ (+1.5) 38.3∗∗∗ (–3.2) 38.5∗∗∗ (–3.0) 40.4∗∗∗ (–1.1) 29.1∗∗∗ (–12.4)
qwen3 28.2∗∗∗ 29.3∗∗∗ (+1.1) 27.3∗∗∗ (–0.9) 26.9∗∗∗ (–1.4) 26.5∗∗∗ (–1.7) 14.9∗∗∗ (–13.4)
Note: $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$
Robustness
We evaluated 18 different LLMs on this benchmark, and results are summarized in Table 1. For each variation of the model, we also conducted a two-proportion z-test to measure if the accuracy rate decreases significantly as compared to the original. Statistically significant differences are indicated using standard notation (*$p<0.1$, **$p<0.05$, ***$p<0.01$). We also computed 95 % CI (See Appendix E Figure 4) and robustness metrics $R$ (See Appendix E Figure 7), and all models, especially those performed well on the original set, got a low robustness score.
We observe that almost all variants lead to a decrease in model accuracy, even when the transformation is merely changing the names of the variables. This indicates a notable lack of robustness: models often lack the capability to preserve their accuracy under mathematically identical but surface-modified representations. Particularly, transformations that rely on variable-name reasoning (such as Misleading or Garbled String) tend to disturb the model’s math accuracy most severely.
Another observation is if a model is not robust on one variant, it tends to be not robust on other variants. Notable examples would be kimi-k2, cluase-opus-4 and gemini-2.5-pro.
Transformation-wise Breakdown
We now present a detailed comparison of model robustness across the five variant types: Descriptive, Confusing, Misleading, Garbled String, and Kernel Variant.
Descriptive Long (DL)
This transformation had the smallest impact on accuracy across all models. The average accuracy drop is marginal, and most drops are not statistically significant. Some models such as o3 (+1.3), o4-mini(+1.5), and Qwen3-235B (+1.1) even improved slightly. As expected, changing variables to descriptive names tends to preserve accuracy.
Confusing (DLC)
The addition of semantically meaningless but structurally longer variable names caused a moderate decrease in accuracy. Models like Claude-opus-4 (–4.3**) and GPT-4o-mini (–1.8**) showed statistically significant drops.
Misleading (DLM)
Perturbing the questions by replacing variable names with misleading strings has a strong impact on mathematical reasoning accuracy. Nearly all models experienced a significant drop, with several exceeding –5 percentage points. Notably, Claude-Opus-4 (–4.8**), Gemini-2.5-pro (–5.6***), and Claude-Sonnet-4 (–4.4**) were among the most heavily affected.
Garbled String (GS)
Replacing variable names with random character strings consistently degraded performance across all LLMs. Every model lost accuracy, and more than half of them had drops that reached statistical significance. Models such as Gemini-2.5-pro (–5.4***), Claude-Sonnet-4 (–4.9***), and Gemini-2.5-flash-lite (–4.7***) suffered the largest declines.
Kernel Variant (KV)
Kernel variants—which keep each question’s mathematical structure but replace constants and expressions with different values—led to the sharpest decline overall. All models experienced large drops, often in the range of –5 to –15 points, with Grok4 (–14.6***), Gemini-2.5-flash (–15.2***), and Gemini-2.5-pro (–14.9***) showing the steepest declines.
Overall, these results highlight that even state-of-the-art LLMs struggle to maintain consistent performance under simple, semantics-preserving transformations. These findings suggest that LLMs may not consistently invoke their mathematical reasoning abilities, even when such knowledge is likely embedded in their parameters. Instead, performance appears sensitive to superficial cues such as variable names or textual formatting—features that should be irrelevant to a mathematically robust solution.
Error Taxonomy
Our grading script returns a brief comment for every incorrect answer. Using these comments, we grouped errors into four categories: Symbol Confusion, Step Omission, Arithmetic, and Logic Hallucination. Figure 3 shows that the relative frequency of these error types is nearly identical across variants; logic hallucinations dominate, accounting for roughly three-fifths of all wrong answers regardless of prompt wording. Thus, the accuracy drop is distributed across all categories rather than driven by a single one, confirming that mathematically equivalent perturbation consistently degrades LLM performance.
Figure 3: Error Composition Ratio across Variants
Discussion
Key Findings
Symbol-level perturbations cause substantial drops.
Across the four surface variants—DL, DLC, DLM, and GS—merely renaming variables lowers accuracy by $3\text{--}5$ pp on average; for example, Gemini-2.5-pro falls from 78.3% to 72.9% (–5.4 pp; see Table 1). This indicates that today’s SOTA models still rely on lexical “semantic anchors” rather than fully abstract proof structures.
Maintaining structure but resampling parameters is even harsher.
The Kernel Variant (KV) simultaneously resamples all mutable constants while preserving the original reasoning skeleton. Accuracy losses reach $\approx 10$ pp; OpenAI O3 declines from 48.8% to 38.5% (–10.3 pp), showing that grasping a solution pattern does not automatically translate to parameter-invariant reasoning ability.
$R_{\text{global}}$ reveals fine-grained brittleness.
We compute $R_{\text{surf}},R_{\text{core}},R_{\text{global}}$ where $R(\cdot,\cdot)$ is the SD–normalized robustness metric. Because it exponentially penalizes rare but catastrophic flips, $R_{\text{global}}$ tracks effective robustness more faithfully than a plain hard/easy accuracy ratio.
Implications
Evaluation. The GAP framework generates an (in principle) unbounded supply of semantically equivalent test items, limiting future benchmark leakage and mitigating leaderboard inflation.
Training. Results suggest curriculum fine-tuning that explicitly randomizes (i) symbol identities and (ii) numeric parameters, instead of simply enlarging pre-training corpora.
Security. Surface-level fragility implies that production systems can be prompt-injected with mathematically innocuous renamings—highlighting the need to integrate robustness checks into red-team pipelines.
Practical Recommendations
Our study suggests that some strategies such as the following may potentially inprove the performance of LLMs on math reasoning tasks.
  1. 1.
    Data augmentation. Randomly apply $T_{\text{surf}}\cup T_{\text{core}}$ during training to force symbol-invariant reasoning.
  2. 2.
    Symbol binding. Separate identifier tokens from literal tokens (e.g., via a learnable symbol table) inside the Transformer.
  3. 3.
    Hybrid reasoning. Embed SMT/CAS validators into decoding (e.g., value-head alignment) to tighten logical consistency.
Ethical Considerations & Societal Impact
Our benchmark is released under a non-commercial licence with variants and auto-graders only; raw solutions remain withheld. This transparency enables reproducible stress tests while limiting the risk of seeding training corpora with answer keys. Nonetheless, the same techniques could craft adversarial prompts that mislead automated theorem provers, so we encourage multi-agent verification in high-stakes deployments.
Limitations
LLM‑generated variants and verifier bias.
The O3‑based generator and five‑judge verifier bound the chance of accepting a flawed core‑step variant to $\delta<10^{-10}$. Because all judges share the same architecture, rare shared oversights may persist; future versions will mix in symbolic provers or heterogeneously trained LLMs to strengthen the safety net.
Automatic‑grading noise.
Our LLM rubric delivers $\approx 97\%$ precision/recall, leaving a uniform $\lesssim 3\%$ label error that has so far not altered model rankings. Multi‑agent grading or targeted human audits on disagreement cases could tighten this bound.
Residual data leakage.
Kernel resampling and systematic symbol renaming greatly reduce verbatim memorization, yet structurally similar proofs in pre‑training data remain a remote possibility. Hash‑challenge prompts or evaluation‑time watermarking would help detect such leakage.
Metric interpretability.
The SD‑normalized robustness score $R$ is scale‑free and differentiable, but its exponential compression can mask item‑level failure modes. We therefore pair $R$ with distributional diagnostics such as Batched Drop Curves to improve transparency.
Related Work
There have been multiple benchmarks for evaluating the mathematical-reasoning capabilities of large language models (LLMs). Early math‐reasoning benchmarks such as MATH(1.25 k problems), and GSM8K(8.5 k problems), revealed basic arithmetic/algebra skills. (Hendrycks et al. 2021; Cobbe et al. 2021). But their difficulty is now saturated as LLMs scale. For instance, with prompting strategies such as DUP, GPT-4 attains 97.1% accuracy on GSM8K (Zhong et al. 2025). This ceiling at the high-school-competition level motivated the creation of a new generation of harder benchmarks.
Subsequent benchmarks target harder problems. Omni-MATH contributes 4 428 rigorously annotated Olympiad-level problems (Gao et al. 2024). Likewise, OlympiadBench provides a bilingual, multimodal benchmark of 8 476 Olympiad-level math and physics problems with expert step-by-step solutions (He et al. 2024). The cross-disciplinary benchmark ARB consist questions in mathematics, physics, biology, chemistry, and law, with a rubric-based self-grading protocol (Sawada et al. 2023). Some other benchmarks focuses specifically on formal proof. MiniF2F supplies 488 Olympiad-level problems formalized in multiple proof assistants (Zheng, Han, and Polu 2022). Putnambench, offers 1 692 rigorously hand-crafted formalizations of Putnam Competition problems(Tsoukalas et al. 2024).
Nevertheless, recent studies warn that scores on many NLP benchmarks may be artificially inflated by data contamination, when LLMs are trained on the benchmark questions. Sainz et al. point out that many benchmarks may be inflated because large language models often memorize test data seen during pre-training (Sainz et al. 2023). Balloccu et al. conduct a systematic audit of data leakage for closed-source LLMs and estimate that roughly 4.7 million test examples from 263 datasets were likely exposed to the models (Balloccu et al. 2024).
In order to obtain a more robust evaluation of LLMs’ reasoning capabilities, it is important to prevent data leakage. One approach is to create original questions. FrontierMath, for example, addresses this with a rigorously curated benchmark of hundreds of original, expert-level mathematics problems spanning fields from number theory to algebraic geometry (Glazer et al. 2024). Putnam-AXIOM takes this approach and comprises 522 challenging problems from the William Lowell Putnam Competition plus 100 programmatically generated functional variations, furnishing a contamination-resilient benchmark (Gulati et al. 2025).
Another method to deal with data leakage is introducing contrast sets-small, label-changing perturbations of existing test instances—to probe a model’s local decision boundary (Gardner et al. 2020). Huang et al. construct MATH-Perturb that applies simple and hard perturbations to 279 level-5 MATH problems and discover that models suffer 12–16 percentage-point drops on the hard variant(Huang et al. 2025).Shalyt et al. complement this line of work with ASyMOB, a 17 k-problem benchmark whose algebra-focused numerical and symbolic perturbations reveal performance drops of up to 70 percentage points, highlighting models’ fragility under such stress tests (Shalyt, Elimelech, and Kaminer 2025). Similarly, Yu et al. propose Math-RoB, a synthetic benchmark that enables robustness evaluation against the data contamination through an instruction-based approach(Yu et al. 2025). These efforts either focus on specific aspects that limit generalizability, are based on benchmarks that are too easy for current models, or introduce transformations that are not mathematically equivalent, thus confounding true robustness evaluation.
Building on these prior efforts, our work introduces Generalization–and–Perturbation (GAP), a unified framework that addresses both data leakage and robustness by generating mathematically equivalent variants of complex problems, significantly expanding the evaluation depth of existing benchmarks. The framework can be applied to existing and future benchmarks, and all types of questions to strengthen their reliability. To address the saturation of accuracy scores, we apply our framework to challenging, college-level competition mathematics problems. We instantiate GAP on every William Lowell Putnam Competition problem from 1938–2024 (1 051 originals), expanding each item into five mathematically equivalent variants and thereby producing PutnamGAP, a corpus of 6,306 stress-test questions. Finally, we release an open-source evaluation stack that rigorously grades solutions step by step, making assessment fully automated, transparent, and reproducible.
Conclusion & Future Work
In this paper, we introduced the Generalization–and–Perturbation (GAP) framework. By instantiating GAP on all 1 051 Putnam Competition questions we produced the 6 306‑question PutnamGAP benchmark. A zero‑shot evaluation of 18 commercial and open‑source LLMs revealed sharp and consistent accuracy drops. These results expose a clear robustness gap that leaderboard scores on unperturbed datasets have so far not shown.
The findings highlight three actionable directions.
  • Benchmarking: GAP offers an open‑ended supply of contamination‑resistant test items, limiting future data leakage and score inflation.
  • Training: curricula that randomize both symbol identities and numeric parameters during fine‑tuning should become standard practice for models targeting formal reasoning domains.
  • Security: the same surface‑level fragility that hurts accuracy can be weaponized for prompt‑injection attacks, so GAP‑style mutation should be built into red‑teaming pipelines.
Moving forward we will (i) diversify the verifier ensemble with symbolic provers and heterogeneous LLMs to rule out collusive blind spots, (ii) port GAP to applied mathematics, physics and multi-modal STEM corpora, and (iii) integrate on‑the‑fly GAP transformations into training so that invariance to symbol and parameter changes is learned rather than merely tested.
PutnamGAP makes one lesson unmistakable: genuine progress in mathematical AI will be measured not by ever‑higher raw scores, but by a model’s ability to stride across the hidden gulf between symbols and substance. The next generation of top‑tier systems will earn their place only by refusing to be left behind on GAPs.
Appendix A
Model Params (B) Availability
GPT‑4.1‑nano Undiscl. API
GPT‑4.1‑mini Undiscl. API
GPT‑4.1 Undiscl. API
GPT‑4o‑mini $\sim$8 API
GPT‑4o $\sim$1 800 (MoE) API
o3 Undiscl. API
o4‑mini Undiscl. API
Gemini-2.5-flash‑lite Undiscl. API
Gemini-2.5-flash Undiscl. API
Gemini-2.5-pro $\gg$100* API
Claude-Sonnet-4 Undiscl. API
Claude-Opus-4 Undiscl. API
Grok4 Undiscl. API
Kimi-K2 1 000 / 32 (MoE) Open Source + API
Llama4 Maverick 400 / 17 (MoE) Open Source + API
DeepSeek-Prover-V2 671 / 37 (MoE) Open Source + API
Qwen-3 235B 235 / 22 (MoE) Open Source + API
Mistral Devstral‑Medium Undiscl. API
Table 2: Models evaluated in this work. = built‑in two‑stage “reasoning” pipeline. MoE rows list total / per‑token activated parameters. * = community estimate.
Appendix B
Motivation.
Benchmark leakage inflates raw accuracy; what matters is how much a hard re‑phrasing degrades performance relative to an easy variant of the same problem. A useful metric must be (i) item‑aware (catastrophic flips hurt more than mild drops), (ii) scale‑free (comparable across tasks and models), and (iii) differentiable should one wish to optimize for robustness. The following definition satisfies all three.  [^1] We later introduce a softplus saturation to avoid unbounded penalties for very rare but extreme flips while preserving differentiability.
Notation
Let an LLM’s binary correctness on $N$ aligned items be
$\mathbf{e}=(e_{1},\dots,e_{N}),\qquad\mathbf{h}=(h_{1},\dots,h_{N})\in\{0,1\}^{N},$
where $\mathbf{e}$ refers to the easy (original) statements and $\mathbf{h}$ to their hard counterparts. We later extend $e_{j},h_{j}$ to lie in $[0,1]$ to accommodate soft probabilities; in the binary case they take values in $\{0,1\}$. Jeffreys smoothing prevents pathologies when a model is perfect (or useless) on one split:
$p_{e}=\frac{\sum_{j}e_{j}+\nicefrac{{1}}{{2}}}{N+1},\qquad p_{h}=\frac{\sum_{j}h_{j}+\nicefrac{{1}}{{2}}}{N+1}.$ (4)
The pooled standard deviation is
$\sigma=\sqrt{\tfrac{1}{2}\!\bigl{[}p_{e}(1-p_{e})+p_{h}(1-p_{h})\bigr{]}}.$ (5)
SD‑normalized drop
For every aligned pair we express the change in SD units:
$d_{j}=\frac{e_{j}-h_{j}}{\sigma}\;\in\;\{-1,0,1\}/\sigma.$ (6)
A positive $d_{j}$ means the hard item flipped to wrong; a negative $d_{j}$ means the model solved the hard variant only.
$\hat{d}_{j}=\frac{1}{k}\ln\!\bigl{(}1+e^{k\,d_{j}}\bigr{)},\qquad k\approx 0.5.$ (7)
For $k\to\infty$ the soft‑plus collapses to $d_{j}$.
Data‑driven scaling
Let $\tilde{d}=\operatorname*{median}\{d_{j}\,\mid\,d_{j}>0\}$ denote the median positive drop. We choose an exponential mapping whose slope $\beta=\ln 2/\tilde{d}$ forces this “typical” loss to score $\nicefrac{{1}}{{2}}$. When no positive drop exists, we adopt the data‑driven fallback
$\tilde{d}:=\max\!\bigl{(}\varepsilon,\;\operatorname*{median}|d_{j}|\bigr{)},\qquad\varepsilon=0.1.$
Per‑item reward and aggregate robustness
The per‑item score is $r_{j}=\exp(-\beta\,d_{j})$. Finally, the aggregate robustness of model $f$ is
$\boxed{R(\mathbf{e},\mathbf{h})=\frac{1}{N}\sum_{j=1}^{N}\exp\!\Bigl{(}-\tfrac{\ln 2}{\tilde{d}}\,d_{j}\Bigr{)}}$ (8)
A score of 1 means no net change; a one‑SD loss reduces $r_{j}$ to $\approx 0.71$; a two‑SD loss halves it; symmetric gains push $R>1$.
$\boxed{\hat{R}(\mathbf{e},\mathbf{h})=\frac{1}{N}\sum_{j=1}^{N}\exp\!\Bigl{(}-\tfrac{\ln 2}{\tilde{d}}\,\hat{d}_{j}\Bigr{)}}$ (9)
Equation (9) is strictly differentiable for all $e_{j},h_{j}\in[0,1]$ thanks to (7).
Why not the hard/easy accuracy ratio?
Table 3: Key differences between the naive ratio $A_{h}/A_{e}$ and the proposed robustness $R(\mathbf{e},\mathbf{h})$.
Aspect Ratio $A_{h}/A_{e}$ Robustness $R$ (ours)
Granularity Single fraction ignores which items flipped Per‑item drops aggregated; catastrophic flips dominate
Baseline sensitivity Undefined/unstable if $A_{e}\!\to\!0$ Jeffreys smoothing makes $R$ well‑defined everywhere
Rewarding improvements Collapses to >1>1 with no scale Smoothly increases; one‑SD gain yields $r_{j}\!\approx\!1.41$
Penalizing severe drops Linear; many tiny drops $\approx$ few huge ones Exponential, convex; a few large $d_{j}$ hit harder
Cross‑task comparability Not scale‑free $d_{j}$ in SD units; $\beta$ auto‑scaled → comparable
Optimizer friendliness Piece‑wise, flat for correct pairs Log‑soft‑saturated, fully differentiable
Interpretation.
Equation (8) has no tunable hyper‑parameters beyond the statistical constant $\ln 2$; everything else adapts to the empirical difficulty of the model–dataset pair. We report $R$ alongside the raw easy/hard accuracies to keep its meaning transparent. [^2] When $A_{h}>A_{e}$, $R$ can exceed 1; such cases occur only for a small fraction of models/items but deserve inspection.
Connection to Cohen’s $d$
Averaging the per‑item normalised drops recovers the classical effect size:
$\frac{1}{N}\sum_{j=1}^{N}d_{j}\;=\;\frac{p_{e}-p_{h}}{\sqrt{\tfrac{1}{2}[p_{e}(1-p_{e})+p_{h}(1-p_{h})]}}\;=\;d_{\text{Cohen}}.$
Thus our robustness metric can be viewed as an exponential re‑scoring of the sample‑wise components of Cohen’s $d$.
Soft‑probability variant
When $e_{j},h_{j}\in[0,1]$ represent correctness probabilities obtained from the model’s logits, all preceding formulas remain valid. The gradient w.r.t. a soft label $e_{j}$ is
$\frac{\partial\hat{R}}{\partial e_{j}}=\frac{1}{N}\exp\!\Bigl{(}-\tfrac{\ln 2}{\tilde{d}}\,\hat{d}_{j}\Bigr{)}\left[-\frac{\ln 2}{\tilde{d}\sigma}\right]\bigl{(}1-\sigma^{-2}(e_{j}-h_{j})/2\bigr{)},$
which is continuous and suitable for back‑propagation.
Appendix C
To disentangle symbol sensitivity from reasoning transfer, we create two orthogonal families of meaning‑preserving variants for each canonical item $x_{i}$ (Section LABEL:sec:problem-formulation). Surface variants alter only the var / param strings, whereas core‑step variants resample numerical constants while enforcing the original logical skeleton.
Surface Variants
We probe symbol‑level generalisation by automatically renaming every var or param token extracted during pre‑processing, while keeping all scientific constants (sci_const) fixed. A single call to o3 proposes a replacement conditioned on the token role (“free variable” vs. “fixed parameter”), and a post‑validation step rejects any collision with existing identifiers.
For each original problem we synthesise four independent renaming families and instantiate exactly one variant per family, yielding in total $1\,051\times 4=4\,204$ surface items. The families are:
  1. 1.
    Descriptive‑Long (DL). A single, meaningful English phrase (e.g. populationDensity). Accuracy on DL is empirically indistinguishable from the original and therefore serves as a sanity check.
  2. 2.
    Descriptive‑Long‑Confusing (DLC). A concatenation of 2–5 unrelated words (e.g. walnutVioletTerrace), designed to overload working memory without changing semantics.
  3. 3.
    Descriptive‑Long‑Misleading (DLM). A phrase built from mathematical jargon that suggests a different concept—e.g. primeFieldOrder used as a real variable—to test whether models latch onto spurious lexical cues.
  4. 4.
    Garbled‑String (GS). A 4–16 character alphanumeric hash (e.g. xcQ7h2ZfRw9v), eliminating any linguistic hint.
Core‑step Variants
While surface renaming stresses symbol recognition, we also wish to test whether a language model can transfer the reasoning skeleton to a numerically distinct yet logically equivalent instance. For every original item we therefore generate a single core‑step variant via the four‑stage pipeline:
  1. 1.
    Slot discovery Forward $(x_{i},\pi_{i})$ to o3; it lists every constant whose value is not logically fixed, emitting a mutable_slot dictionary with human‑readable descriptors (e.g. “neighborhood half‑width DD”).
  2. 2.
    Safe resampling. Each slot is resampled uniformly within a guard range derived from the problem’s own inequalities, yielding $\{\widetilde{D},\widetilde{k}$$\dots\}$.
  3. 3.
    Back‑synthesis We feed $\langle x_{i},\texttt{slots},\pi_{i},\texttt{mutable\_steps}\rangle$ back to o3; it fills the new constants and regenerates a proof whose step order matches mutable_steps, along with the fully worded problem statement.
  4. 4.
    Unanimous judging Five o3 judge instances, each with an independent temperature seed, must all return “solvable and correct”. A rejection auto-triggers patching and re‑verification. After three consecutive clean passes we perform a 10% human audit.
The output artifact, denoted kernel_variant, stores the new statement, regenerated proof, slot dictionary, and preserved core‑step list. Exactly one kernel variant is produced per source item, totaling 1 0511\,051 items.
Theoretical Guarantees
The variant pipeline combines stochastic LLM generation with a repair‑and‑verify loop (Algorithm 1). Although 76.4 % of the corpus are proof‑based items—i.e. cannot be validated by simple numeric inequalities—we prove that the acceptance criterion yields an exponential safety margin.
Notation
Each candidate undergoes at most $T=15$ verification iterations. Within one iteration $t$ we launch $J=5$ independent o3 judges, each returning accept (1 bit) or reject. Denote by $\varepsilon=\Pr[\text{judge mis‑accepts a flawed candidate}]$. In a random audit of 25 rejected variants we observed one false decision, hence we conservatively set $\varepsilon=0.04$.
An iteration $t$ is passed when all $J$ judges vote accept. A candidate is accepted by the pipeline if it passes in two consecutive iterations; otherwise the loop either repairs the artifact or aborts after 15 attempts. A 10% manual audit follows.
$\delta$‑Soundness under two‑in‑a‑row rule
Let $K=2$ be the required streak length. Under independent‑judge assumption the probability that an unsolvable or incorrect variant survives the pipeline is bounded by
$\delta\;\leq\;(T-K+1)\;\varepsilon^{\,KJ}\;=\;14\;\varepsilon^{10}\;\approx\;14\times(0.04)^{10}<10^{-10}.$
The pipeline examines at most TK+1=14T\!-\!K\!+\!1=14 distinct length‑$K$ windows $\langle t,\dots,t+K-1\rangle$. For a flawed candidate to be accepted, every judge in both iterations of some window must err, an event of probability $\varepsilon^{KJ}$. A union bound over all windows yields the claim.
Why not pre‑computed guard ranges?
Because the majority (76.4 %) of items require multi‑step proofs, the notion of “feasible numeric interval” is ill‑defined. We therefore rely on the rejection‑sampling loop in Algorithm 1; Theorem $\delta$‑Soundness under two‑in‑a‑row rule shows that its soundness is already more stringent than $10^{-9}$, rendering an extra symbolic guard unnecessary.
Reasoning‑step isomorphism
Stage 3 forces the regenerated proof to match the abstract skeleton mutable_steps step‑by‑step, hence every accepted core‑step variant is isomorphic to the source solution $\pi_{i}$ under the identifier mapping introduced in Section Core‑step Variants. A regex verifier found zero mismatches over all 1 051 core variants.
Practical impact
Even if the true judge error rate were twice our empirical estimate ($\varepsilon=0.08$), the bound remains $\delta<10^{-8}$. Thus all reported robustness numbers are statistically safe from false positives introduced by the generation machinery.
Appendix D
We obtain every official problem of the William Lowell Putnam Mathematical Competition from 1938 to 2024 by digitizing the four authoritative monographs shown in Table 4. Each volume is issued by the Mathematical Association of America (MAA) and reprinted by the American Mathematical Society (AMS) under the MAA Press Problem Books series. [^3] Softcover and e-book reprints are available from https://bookstore.ams.org.
Volume (Years) Reference
I (1938–1964) Gleason, Greenwood, and Kelly (1980)
II (1965–1984) Alexanderson, Klosinski, and Larson (1985)
III (1985–2000) Kedlaya, Poonen, and Vakil (2002)
IV (2001–2016) Kedlaya et al. (2020)
Table 4: Primary sources for PutnamGAP. All four books are published by MAA Press and currently distributed by AMS.
Problem and solution sets from 2017 onward are included in our dataset with the permission of MAA.
Across the early era (1938–1941) the competition featured 6688 problems per part (A and B); from 1942 onward the format stabilised at 5566 problems per part, with difficulty increasing monotonically from position 1 to 6.444A few years, such as the wartime years 1943–1945, were canceled; our index skips these years. These historical variations are preserved in our metadata and later support the difficulty-gradient analysis in section Statistics
Extraction & Annotation Pipeline
Our raw sources are scanned PDFs; no machine‑readable  is provided. We therefore build a four‑stage pipeline that converts each page into a fully annotated problem record suitable for variant generation and automatic scoring.
1. Image segmentation & OCR.
Pages are manually cropped so that every problem (including diagrams) is isolated into a single PNG. We then send the image to MathPix, receiving that compiles without error. Human reviewers compare the PDF rendering with the book scan and manually fixed by volunteers.
2. Minimal  normalisation.
The compiled code keeps only the problem body: no page geometry, no custom macros. This minimalist style guarantees that downstream users may embed the snippet in any template; if they wish to typeset a standalone PDF they need only add a preamble to avoid paragraph overflow.
3. Semantic annotation via LLM
Given the cleaned “problem + solution” pair, we prompt OpenAI’s o3 model to extract three kinds of metadata:
  1. 1.
    Topical tags drawn from problem categories $\{\mathrm{ALG},\mathrm{NT},\mathrm{COMB},\mathrm{GEO},\mathrm{ANA}\}$. The tag most central to the pivotal lemma is stored as the unique type. These tags allow users to filter, e.g. “geometry only” subsets.
  2. 2.
    Symbol inventory $\{\texttt{var},\texttt{param},\texttt{sci\_const}\}$: var denotes free variables, param denotes numeric parameters fixed in the statement, and sci_const collects immutable objects like $\pi$ or $e$. During surface‑variant generation we replace only var/param so that scientific constants remain intact.
  3. 3.
    Mutable step list $(S_{1},\dots,S_{k})$. We first linearly segment the official solution into atomic sentences. The cleaned problem–solution pair is then fed to o3, which returns for every sentence the mathematical operation performed (e.g. “integrate by parts”, “apply AM–GM”). We keep the resulting ordered list as core_steps, storing it verbatim in _meta in our dataset. These machine‑readable steps later serve two purposes: (a) guiding the core‑step variant generator, which rewrites the narrative while preserving the entire chain of reasoning, and (b) enabling fine‑grained error analysis of which techniques an LLM fails to execute.
Appendix E
Figure 4: Accuracies of each variant per model bar plot
Figure 5: Accuracies of five types of questions per model per variant heatmap
Figure 6: Accuracies of five types of questions for each variant per model
Figure 7: Accuracies of two classes of questions for each variant per model
Figure 8: Robustness $R$ of questions for each variant per model
Appendix F
Other observations
Some reasoning models get into dead loops during reasoning process until reaching the time limit, making the benchmark users have no choice but to run the tests again to avoid lowering their score due to such time limits, potentially changing PASS@1 into PASS@K and improving the performance during tests. Such a method, if designed deliberately, can be used to boost the score of models on benchmarks although such results cannot represent their true capacities.