### v2

# An Investigation of Robustness of LLMs in Mathematical Reasoning: Benchmarking with Mathematically-Equivalent Transformation of Advanced Mathematical Problems

Abstract
In this paper, we introduce a systematic framework beyond conventional method to assess LLMs’ mathematical-reasoning robustness by stress-testing them on advanced math problems that are mathematically equivalent but with linguistic and parametric variation. These transformations allow us to measure the sensitivity of LLMs to non-mathematical perturbations, thereby enabling a more accurate evaluation of their mathematical reasoning capabilities. Using this new evaluation methodology, we created PutnamGAP, a new benchmark dataset with multiple mathematically-equivalent variations of competition-level math problems. With the new dataset, we evaluate multiple families of representative LLMs and examine their robustness. Across 18 commercial and open-source models we observe sharp performance degradation on the variants. OpenAI’s flagship reasoning model, O3, scores 51.5 % on the originals but drops by 4.7 percentage points on surface-renaming variants, and by 12.9 percentage points on parametric variants, while smaller models fare far worse. Overall, the results show that the proposed new evaluation methodology is effective for deepening our understanding of the robustness of LLMs and generating new insights for further improving their mathematical reasoning capabilities.
1  Introduction
Motivation. Modern AI systems are increasingly entrusted with tasks that hinge on robust reasoning rather than pattern matching. It is thus important to precisely measure an LLM’s reasoning capacity and its ability to generalize beyond memorized textual surface forms. Existing math-reasoning benchmarks, however, exhibit two critical weaknesses: (i) leakage-induced score inflation, since benchmark items rapidly seep into pre-training corpora, and (ii) limited robustness coverage, because today’s datasets are too small or lack controlled transformations that probe true generalization. Addressing these weaknesses is urgent if we aim to benchmark reasoning with the same rigor demanded in safety-critical domains such as healthcare or cybersecurity.
Benchmark inflation through training leakage. Recent studies show that public datasets, including GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021), have leaked into the web‑scale corpora used to pre‑train large language models (LLMs), artificially inflating test‑time accuracy. A leaderboard score therefore no longer guarantees genuine reasoning ability; it may merely reflect memorization of benchmark items or their solutions. Simply releasing yet another dataset postpones the problem: once its items enter future training corpora, scores climb without real progress. What is needed is a systematic method that (i) measures a model’s capacity to generalize beyond verbatim memory and (ii) can generate an unbounded supply of evaluation items, limiting future leakage.
Figure 1: PutnamGAP variants performance relative to the original set
Competition mathematics reveals the next robustness bottleneck. Large language models (LLMs) now surpass 90% accuracy on widely‑used benchmarks such as GSM8K and MATH, prompting claims of “near‑human” numerical reasoning yet still falter on Olympiad‑style or Putnam‑level problems that intertwine multiple domains. Existing Putnam‑derived datasets are too small to expose this gap: Putnam‑AXIOM (236 originals + 52 variations) (Huang et al., 2025), and PutnamBench (640 formalized theorems) (Tsoukalas et al., 2024) remain in the hundreds, and none delivers systematic generalization and perturbations. These facts expose Weakness (i) insufficient scale and Weakness (ii) lack of controlled, systematic transformations in existing evaluations.
Generalization–and–Perturbation (GAP), a novel evaluation strategy. We address both leakage and robustness with a new idea: stress‑test the model on mathematically equivalent versions of the same problem. For a problem $x$ with solution set $S(x)$ and an LLM $f$, robustness is the expected accuracy when $x$ is transformed by a family $\mathcal{T}$ of equivalence‑preserving operators. We partition $\mathcal{T}$ into $\mathcal{T}_{\text{surf}}$ (surface renames that alter symbol salience) and $\mathcal{T}_{\text{para}}$ (kernel rewrites that preserve the same proof steps while changing the scenario and parameters). This GAP framework (i) creates an infinite stream of unseen test items, mitigating future contamination, and (ii) quantifies how far a model can generalize beyond memorized surface forms. It offers a new general diagnostic evaluation methodology for analyzing and quantifying the robustness of an LLM’s mathematical reasoning capacity.
PutnamGAP: instantiating GAP on 85 years of problems. We instantiate GAP on every William Lowell Putnam Competition problem from 1938–2024 (1,051 originals) and expand each item into five variants—four surface renames and one kernel rewrite—obtaining 6,306 stress‑test questions. A two‑stage QA pass—15 rounds of o3 self‑review plus a 10% spot‑check found no substantive errors.
Headline results. Across 18 models, as shown figure 4, all of them suffer from both simple renaming and step-based rewrites. OpenAI’s o3 scores 51.5% on original statements but loses 4.7 pp (9.12%) under surface renames and 12.9 pp (25.22%) under parametric rewrites. These drops confirm that high leaderboard scores can collapse when cosmetic or structural perturbations are applied—precisely the effect that data leakage masks.
Contributions. (1) We propose GAP, a novel general framework for measuring robustness via mathematically equivalent transformations that overcomes two common deficiencies of the current evaluation methods (i.e., data leakage and lack of robustness measures). (2) We release PutnamGAP, the first 6k‑scale competition benchmark that systematically disentangles surface‑level and structural generalization while limiting future leakage. (3) We provide the first comprehensive robustness baseline across seventeen LLMs, plus an open‑source evaluation stack.
2  The Generalization–and–Perturbation (GAP) Framework
2.1 Evaluation Model
We start from a curated set of $N$ canonical items $\mathcal{P}=\{(x_{i},y_{i},\pi_{i})\}_{i=1}^{N}$, where $x_{i}$ is a problem statement, $y_{i}$ is its reference answer(s), and $\pi_{i}$ an unreleased expert solution path used internally for safe variant generation. Model interface. A language model $f_{\theta}$ receives a prompt $x$ and returns $\hat{y}=f_{\theta}(x)$, which an automatic checker maps to a binary label $z=\operatorname{grade}(\hat{y},y)\in\{0,1\}$.
Variant families. For every $x_{i}$ we later apply two disjoint transformation super-families (defined in the next section but left unchanged here): $\mathcal{T}_{i}^{\textsf{surf}}$ ($K_{surf}$ surface variants), $\mathcal{T}_{i}^{\textsf{para}}$ ($K_{para}$ parametric variants). Each surface transformation $\tau$ returns a new statement $x_{i}^{(\tau)}=\tau(x_{i})$ that preserves semantic correctness of $y_{i}$. For parametric variations, $y_{i}$ is transformed as well to match $\tau(x_{i})$.
Evaluation matrix. The Cartesian product $\mathcal{D}=\{(i,\tau)\mid i\leq N,\;\tau\in\mathcal{T}_{i}^{\textsf{surf}}\cup\mathcal{T}_{i}^{\textsf{para}}\cup\{\mathrm{id}\}\}$ contains $N\times(K+1)$ aligned items (original + $K$ variants per source, $K=K_{surf}+K_{para}$). Running $f_{\theta}$ on every pair populates a binary matrix $\mathbf{Z}\in\{0,1\}^{N\times(K+1)}$. From the first column we extract the easy vector $\mathbf{e}(\theta)\in\{0,1\}^{N}$, while the remaining columns feed family-specific aggregates: $\mathbf{h}^{\textsf{surf}}(\theta)=\text{maj}(\mathbf{Z}_{[:,\,\textsf{surf}]})$, $\mathbf{h}^{\textsf{para}}(\theta)=\mathbf{Z}_{[:,\,\textsf{para}]}\!$. The set of surface variants can be changed based on specific tasks.
Robustness Metric. Let $e,h\in\{0,1\}^{N}$ denote per-item correctness on the easy (original) and hard (variant) sets. With Jeffreys smoothing
$p_{e}=\frac{\sum_{j}e_{j}+\tfrac{1}{2}}{N+1},\quad p_{h}=\frac{\sum_{j}h_{j}+\tfrac{1}{2}}{N+1},\quad\sigma=\sqrt{\tfrac{1}{2}\big(p_{e}(1-p_{e})+p_{h}(1-p_{h})\big)}.$
Define the SD-normalized drop $d_{j}=(e_{j}-h_{j})/\sigma$ and its soft-saturated version $\widehat{d}_{j}=\tfrac{1}{k}\log\!\big(1+e^{kd_{j}}\big)$ with $k\approx 0.5$. Let $\tilde{d}=\operatorname{median}\{d_{j}\,|\,d_{j}>0\}$ (with fallback $\tilde{d}:=\max(\varepsilon,\operatorname{median}|d_{j}|)$, $\varepsilon=0.1$ when no positive drop exists) and set $\beta=\ln 2/\tilde{d}$. Our penalty robustness is
$\widehat{R}(e,h)=\frac{1}{N}\sum_{j=1}^{N}\exp\!\big(-\beta\,\widehat{d}_{j}\big)\in(0,1].$
Thus $\widehat{R}=1$ indicates invariance; a “typical” loss ($\widehat{d}_{j}\approx\tilde{d}$) halves the per-item factor, while improvements ($d_{j}<0$) are clamped to zero penalty (no reward). We report $R_{\text{surf}}=\widehat{R}(e,h_{\text{surf}})$, $R_{\text{para}}=\widehat{R}(e,h_{\text{para}})$, and $R_{\text{global}}=\sqrt{R_{\text{surf}}R_{\text{para}}}$. Full derivation, statistical justification, and design discussion are in Appendix B.
2.2 Transformation Families
The proposed general robustness measures can work for any variations. As a first step in exploring this new evaluation methodology, we propose and study five aligned variants— four surface renamings that perturb only symbol names, and one core‑step instance that perturbs numeric slots while preserving the reasoning chain. This section details the synthesis pipelines. Detailed descriptions can also be found in Appendix A.
2.2.1 Surface renaming variant family
Figure 2: Surface renaming variant family pipeline
We want to know whether a model recognizes an argument because it has truly abstracted the pattern or merely because it memorizes suggestive identifier strings. Therefore we systematically replace each token tagged var or param; all constants of category sci_const remain untouched.
Automated pipeline.
  1. 1.
    Proposal. A single call to o3 receives the token role (“free variable” or “fixed parameter”) and the surrounding textual context, and returns a candidate replacement.
  2. 2.
    Collision check. A deterministic post-validator rejects names colliding with any pre-existing identifier in the problem.
  3. 3.
    Family tagging. The string is labelled as belonging to one of four families described below.
We use four types of surface variants: Descriptive_Long (DL), with a single descriptive phrase; Descriptive_Long_Confusing (DLC), with 2–5 random unrelated nouns; Descriptive_Long_Misleading (DLM), with a mathematically suggestive but misleading term ; Garbled_String (GS), with a 4–16-character hash, as shown in figure 2 where ‘Q’ stands for the problem question and ‘A’ stands for the official solution.
Each source item thus yields $4$ surface variants; accuracy deltas per family appear in Section Results & Analysis.
2.2.2 Parametric variant family
Figure 3: Parametric variant family pipeline
Symbol renaming probes only the lexical axis. To probe structural transfer, we resample numerical constants yet force the solution to reuse the original high-level moves. In this work, we call it Kernel_Variant (KV). We convert each item into semantically-equivalent variants through a four-stage pipeline: (1) slot discovery; (2) template back-synthesis; (3) question reverse-engineering; and (4) dual-verifier screening (two-in-a-row rule). The pipeline generates a bounded number of validated variants for each problem within a few hours on commodity hardware using the OpenAI o3 API. See Appendix A for empirical bounds and details of our implementation.
2.3 Implementation Overview
Code release. To facilitate double‑blind reviewing we publish only the subset of data (100 randomly chosen examples). An automated evaluator, putnam-cli.py, receives the names of target solver model and grader model and variant type to test. Supported back‑ends are (i) any HuggingFace‐compatible checkpoint via transformers, (ii) a local vllm server, or (iii) API clients including OpenAI, Gemini, Anthropic and OpenRouter. Full data and generation scripts will be released post‑decision.
Surface generation. Renaming variants are produced on a CPU‑only node by streaming o3 API calls. A five‑stage exponential‑back‑off retry (max 5 attempts, doubling timeout each time) masks transient API latency. Processing all 1 0511\,051 items in parallel takes $\sim$15 min wall‑clock.
Core‑step generation. Kernel variant synthesis is more expensive because of multi‑turn chain‑of‑thought reasoning: end‑to‑end runtime is $\leq\!3\,\text{h}$ for the full corpus on a single 8‑core CPU, dominated by the 15‑iteration repair‑and‑verify loop.
3  PutnamGAP Dataset
3.1 Data Sources, Extraction & Annotation
Our benchmark comprises all Putnam Problems 1938–2024 (N=1 051N=1\,051 items after deduplication). See Appendix E for source details.
Original scans are processed via a 3-stage OCR routine: (i) Manual segmentation for every question-answer pair. (ii) MathPix for formula‐aware PDF-to-LaTeX conversion followed by (iii) custom post-filters that merge multi-line expressions and fix 4̃.2 % residual symbol errors. Each item is manually spot-checked ($\leq$2 min per problem) to ensure semantic fidelity before variant generation. Complete corpus list, OCR accuracy study, and cleaning scripts appear in Appendix E.
3.2 Dataset Statistics
Overall scale and balance. The benchmark comprises 1,051 original Putnam problems from 1938–2024 and five mathematically equivalent transformations, yielding 6,306 items. Part distribution is balanced (527 A vs. 524 B), and the canonical identifier $\langle\textit{year},\textit{part}\{A,B\},\textit{index}\rangle$ provides a difficulty proxy. Using indices 121\!\!-\!2 as Easy, 343\!\!-\!4 as Medium, and 565\!\!-\!6 as Hard, the corpus contains 32.3 % Easy, 32.3 % Medium, 32.2 % Hard, plus a 3.0 % extra–hard tail (indices 7–8).
Topic coverage and Quality Control Automatic tags in _meta.tag indicate broad mathematical coverage—Algebra (641), Analysis (521), Number Theory (392), Combinatorics (286), and Geometry (239). 803 of the questions are proofs, and 248 of them are calculations. At the same time, every item has undergone single‐pass manual validation.
4  Experimental Setup
The constructed PutnamGAP dataset enables, for the first time, a robust analysis of an LLM’s reasoning capacity. In this section, we describe how we set up the experiments to evaluate the robustness of 18 representative models.
4.1 Model Pool & Prompting
We evaluated 18 models (see 1 or Appendix A for a complete list).All models are queried under a unified zero‑shot template. A system instruction designates the model as “an expert mathematician” and asks it to show all work, while the user message embeds the problem. See Appendix G for our full prompt. We fix temperature=0, top_p=1, and max_tokens=32000 or maximum token amount available in case some models have max_tokens maximum smaller than 32000. for every run except OpenAI O-series which require temperature=1. Solutions are then re‑submitted to a second template that grades the answer: a strict proof rubric for proof items and a lenient numeric rubric for calculation items. Both grader prompts require structured JSON output containing a binary grade field plus detailed feedback. Complete prompt code is available in Appendix G
4.2 Scoring & Auto-Grader
We partition tasks into computation and proof categories and evaluate them with distinct graders.
Computation Each candidate answer is normalized (whitespace, units, LaTeX macros) and passed to two scoring paths: (i) a strict string match against the reference solution; (ii) a latent grader—an LLM prompted to return ‘‘CORRECT’’ or ‘‘INCORRECT’’ given the reference answer and a rubric that disallows partial credit. We adopt path (ii) to mitigate formatting artifacts; if the two paths disagree we mark the item for manual audit (¡1% of cases).
Proof We provide the grader with an aligned, step-by-step reference proof and ask it to assign a binary grade plus a natural-language justification. Any skipped logical step or missing citation triggers a fail. A random 10 % sample is double-checked by independent volunteers; grader precision/recall is >>97 %.
5  Results & Analysis
5.1 Robustness
Table 1: Model Accuracy Rates across Categories (Percent Scale)
Model DL ($\Delta$) DLC($\Delta$) DLM ($\Delta$) GS ($\Delta$) Kernel Variant ($\Delta$)
claude-opus-4 23.0∗∗ (–3.5) 22.2∗∗∗ (–4.3) 21.7∗∗∗ (–4.8) 21.4∗∗∗ (–5.1) 13.8∗∗∗ (–12.7)
claude-sonnet-4 20.6∗∗ (–2.5) 19.8∗∗∗ (–3.2) 18.6∗∗∗ (–4.4) 18.1∗∗∗ (–4.9) 11.1∗∗∗ (–11.9)
deepseek-prover 15.2∗∗∗ (–0.3) 14.0∗∗∗ (–1.5) 12.8∗∗ (–2.7) 13.7∗∗ (–1.8) 9.2 ∗∗∗ (–6.3)
gemini-2.5-flash-lite 18.8∗∗∗ (–0.9) 16.1∗∗∗ (–3.7) 15.8∗∗∗ (–4.0) 15.1∗∗∗ (–4.7) 6.6 ∗∗∗ (–13.2)
gemini-2.5-pro 75.2∗∗ (–3.1) 74.3∗∗∗ (–4.1) 72.8∗∗∗ (–5.6) 72.9∗∗∗ (–5.4) 63.5∗∗∗ (–14.9)
gemini-2.5-flash 42.6∗∗∗ (–0.2) 39.0∗∗∗ (–3.8) 40.9∗∗∗ (–1.9) 37.6∗∗∗ (–5.2) 27.6∗∗∗ (–15.2)
gpt-4.1 23.4∗∗∗ (–1.5) 21.2∗∗ (–3.7) 22.0∗∗ (–2.9) 22.6∗∗∗ (–2.3) 14.8∗∗∗ (–10.1)
gpt-4.1-mini 26.9∗∗ (–1.7) 27.1∗∗∗ (–1.5) 24.0∗∗∗ (–4.6) 25.1∗∗ (–3.4) 19.2∗∗∗ (–9.4)
gpt-4.1-nano 8.9 ∗∗∗ (+0.1) 6.4 ∗∗ (–2.4) 7.3 ∗∗ (–1.5) 6.6 ∗∗ (–2.2) 6.8 ∗∗∗ (–2.0)
gpt-4o 6.3 ∗∗∗ (–0.1) 5.3 ∗∗ (–1.1) 4.7 ∗∗ (–1.8) 4.7 ∗∗∗ (–1.8) 3.0 ∗∗∗ (–3.4)
gpt-4o-mini 3.5 ∗∗∗ (–0.8) 2.5 ∗∗∗ (–1.8) 3.3 ∗∗∗ (–1.0) 3.2 ∗∗∗ (–1.1) 1.7 ∗∗∗ (–2.6)
grok4 59.0∗∗∗ (–1.1) 56.6∗∗∗ (–3.4) 55.9∗∗∗ (–4.2) 55.2∗∗∗ (–4.8) 45.5∗∗∗ (–14.6)
kimi-k2 25.8∗∗∗ (–1.4) 23.7∗∗ (–3.5) 23.8∗∗ (–3.4) 23.4∗∗∗ (–3.8) 12.8∗∗∗ (–14.4)
llama-4 14.5∗∗∗ (–1.2) 13.0∗∗ (–2.7) 13.1∗∗ (–2.6) 13.8∗∗ (–1.9) 7.3 ∗∗∗ (–8.4)
mistral 5.5 ∗∗∗ (–0.1) 5.7 ∗∗∗ (+0.1) 4.9 ∗∗∗ (–0.7) 4.2 ∗∗ (–1.3) 3.9 ∗∗ (–1.6)
o3 52.8∗∗∗ (+1.3) 47.6∗∗ (–3.8) 49.5∗∗∗ (–2.0) 46.8∗∗∗ (–4.7) 38.6∗∗∗ (–12.9)
o4-mini 43.0∗∗∗ (+1.5) 38.3∗∗ (–3.2) 38.5∗∗∗ (–3.0) 40.4∗∗∗ (–1.1) 29.1∗∗∗ (–12.4)
qwen3 29.3∗∗∗ (+1.1) 27.3∗∗∗ (–0.9) 26.9∗∗∗ (–1.4) 26.5∗∗∗ (–1.7) 14.9∗∗∗ (–13.4)
Note: $^{***}p<0.01$, $^{**}p<0.05$, $^{*}p<0.1$
We evaluated 18 different LLMs on this benchmark, and results are summarized in Table 1. For each variation of the model, we used a paired design (McNemar’s exact test) on matched problem pairs to test whether the accuracy rate decreases significantly compared to the original. Statistically significant differences are indicated using standard notation ($p<0.1$, $p<0.05$, $p<0.01$). We also computed 95 % CI (See Appendix D Figure 4 ) and robustness metrics $R$ (See Appendix D Figure 9 ), and all models, especially those performed well on the original set, got a low robustness score.
We observe that almost all variants lead to a decrease in model accuracy, even when the transformation is merely changing the names of the variables. This indicates a notable lack of robustness: models often lack the capability to preserve their accuracy under mathematically identical but surface-modified representations. Particularly, transformations that rely on variable-name reasoning (such as Misleading or Garbled String) tend to disturb the model’s math accuracy most severely.
Table 2: Robustness metrics $R_{\text{surf}},R_{\text{para}},R_{\text{global}}$ (rounded to three decimals).
Model $R_{\text{surf}}$ $R_{\text{para}}$ $R_{\text{global}}$ Model $R_{\text{surf}}$ $R_{\text{para}}$ $R_{\text{global}}$
claude-opus-4 0.958 2360.958\,236 0.948 9230.948\,923 0.953 5690.953\,569 gpt-4o 0.985 8360.985\,836 0.980 0260.980\,026 0.982 9270.982\,927
claude-sonnet-4 0.961 3280.961\,328 0.941 5450.941\,545 0.951 3850.951\,385 gpt-4o-mini 0.989 6640.989\,664 0.985 6520.985\,652 0.987 6560.987\,656
deepseek-prover 0.972 3610.972\,361 0.959 8140.959\,814 0.966 0670.966\,067 grok4 0.937 4150.937\,415 0.916 0480.916\,048 0.926 6700.926\,670
gemini-2.5-flash-lite 0.961 2020.961\,202 0.942 2020.942\,202 0.951 6550.951\,655 kimi-k2 0.955 1850.955\,185 0.929 6170.929\,617 0.942 3140.942\,314
gemini-2.5-pro 0.949 1320.949\,132 0.914 8770.914\,877 0.931 8480.931\,848 llama-4 0.972 0190.972\,019 0.954 6040.954\,604 0.963 2720.963\,272
gemini_2.5_flash 0.951 5780.951\,578 0.917 5230.917\,523 0.934 3960.934\,396 mistral 0.984 0810.984\,081 0.981 9830.981\,983 0.983 0320.983\,032
gpt-4.1 0.963 0350.963\,035 0.944 3500.944\,350 0.953 6460.953\,646 o3 0.939 7310.939\,731 0.921 0620.921\,062 0.930 3500.930\,350
gpt-4.1-mini 0.953 1850.953\,185 0.938 5900.938\,590 0.945 8600.945\,860 o4-mini 0.946 0860.946\,086 0.928 9190.928\,919 0.937 4630.937\,463
gpt-4.1-nano 0.980 2450.980\,245 0.982 3470.982\,347 0.981 2950.981\,295 qwen3 0.941 2080.941\,208 0.927 5470.927\,547 0.934 3530.934\,353
[…]
Another observation is if a model is not robust on one variant, it tends to be not robust on other variants. Notable examples would be kimi-k2, cluase-opus-4 and gemini-2.5-pro.
5.2 Transformation-wise Breakdown
Descriptive Long (DL) The impact of this transformation is smallest overall: drops are marginal and mostly not significant. Some models such as o3 (+1.3), o4-mini(+1.5), and Qwen3-235B (+1.1) even improved slightly. This indicates that descriptive renaming preserving accuracy.
Confusing (DLC) Long, semantically meaningless variable names moderately reduce accuracy. Models like Claude-opus-4 (–4.3***) and GPT-4o-mini (–1.8***) showed significant drops.
Misleading (DLM) Replacing variables with misleading strings strongly hurts math accuracy. Nearly all models experienced a significant drop. Notably, Claude-Opus-4 (–4.8***), Gemini-2.5-pro (–5.6***), and Claude-Sonnet-4 (–4.4***) were among the most heavily affected.
Garbled String (GS) Random character strings consistently degrade performance: every model loses accuracy, over half significantly. Models such as Gemini-2.5-pro (–5.4***), Claude-Sonnet-4 (–4.9***), and Gemini-2.5-flash-lite (–4.7***) suffered the largest declines.
Kernel Variant (KV) Kernel variants—which keep each question’s mathematical structure but replace constants and expressions with different values—led to the sharpest decline overall. All models experienced large drops, often in the range of –5 to –15 points, with Grok4 (–14.6***), Gemini-2.5-flash (–15.2***), and Gemini-2.5-pro (–14.9***) showing the steepest declines.
Overall, state-of-the-art LLMs show inconsistent performance under semantics-preserving transformations and appear sensitive to superficial cues. This is consistent with the possibility that part of their gains reflects data-leakage–related memorization rather than stable mathematical reasoning. The pattern persists across topics and problem classes: bar plots with $95\%$ CIs (Appendix D, fig. 4) and per-topic/per-class breakdowns (Appendix D, figs. 7-8) show similar robustness gaps across Algebra/Analysis/NT/Combinatorics/Geometry and for both proof and calculation items.
5.3 Error Taxonomy
Our grading script returns a brief comment for every incorrect answer. Using these comments, we grouped errors into four categories: Symbol Confusion, Step Omission, Arithmetic, and Logic Hallucination. Figure 5 in Appendix D shows that the relative frequency of these error types is nearly identical across variants; logic hallucinations dominate, accounting for roughly three-fifths of all wrong answers regardless of prompt wording. Thus, the accuracy drop is distributed across all categories rather than driven by a single one, confirming that mathematically equivalent perturbation consistently degrades LLM performance.
5.4 External Validation
We applied our surface-renaming protocols—DLC and GS—to ALG514  (kus, 2014). Accuracy decreased from Base 93.6% to DLC 90.9% ($\Delta=-2.7$ pp) and GS 89.3% ($\Delta=-4.3$ pp); McNemar tests (Base vs DLC: b=24,c=10,p=0.024b{=}24,c{=}10,p{=}0.024; Base vs GS: b=35,c=13,p=0.002b{=}35,c{=}13,p{=}0.002). These statistically significant drops indicate that GAP’s surface-renaming stress tests generalize to other math datasets and reveal nontrivial sensitivity to variable renaming.
6  Discussion
6.1 Key Findings
The proposed GAP framework allowed us to make the following new findings about the behavior of LLMs in performing mathematical reasoning:
Symbol-level perturbations cause substantial drops. Across the four surface variants—DL, DLC, DLM, and GS—merely renaming variables lowers accuracy by $3\text{--}5$ pp on average; for example, Gemini-2.5-pro falls from 78.3% to 72.9% (–5.4 pp; see Table 1). This indicates that today’s SOTA models still rely on lexical “semantic anchors” rather than fully abstract proof structures.
Maintaining structure but resampling parameters is even harsher. The Kernel Variant (KV) simultaneously resamples all mutable constants while preserving the original reasoning skeleton. Accuracy losses reach $\approx 10$ pp; OpenAI O3 declines from 48.8% to 38.5% (–10.3 pp), showing that grasping a solution pattern does not automatically translate to parameter-invariant reasoning ability.
$R_{\text{global}}$ reveals fine-grained brittleness. We compute $R_{\text{surf}},R_{\text{para}},R_{\text{global}}$ where $R(\cdot,\cdot)$ is the SD–normalized robustness metric. Because it exponentially penalizes rare but catastrophic flips, $R_{\text{global}}$ tracks effective robustness more faithfully than a plain hard/easy accuracy ratio.
Takeaway. Across capable models we consistently observe $R_{\text{para}}<R_{\text{surf}}$, and we summarize stress-type invariance via $R_{\text{global}}=\sqrt{R_{\text{surf}}R_{\text{para}}}$; interpreting 1R1-R as penalty mass highlights non-trivial fragility even when raw accuracy is high.
6.2 Implications
A novel evaluation methodology: The GAP framework provides a novel methodology for analyzing and evaluating the robustness of LLMs’ reasoning capacity by generating an (in principle) unbounded supply of semantically equivalent test items, which can limit future benchmark leakage and mitigate leaderboard inflation.
Improving robustness via curriculum fine-tuning: Our results suggest curriculum fine-tuning that explicitly randomizes (i) symbol identities and (ii) numeric parameters, instead of simply enlarging pre-training corpora. That is, we can leverage the GAP framework to augment data for fine-tuning a model to improve robustness.
Detecting potential security concerns: Surface-level fragility implies that production systems can be prompt-injected with mathematically innocuous renamings—highlighting the need to integrate robustness checks into red-team pipelines. Our evaluation framework enables such risk analysis before deploying any production system.
Reporting. We recommend reporting bootstrap CIs for $R_{b}$ together with per-item histograms of SD-normalized drops $d_{j}=(e_{j}-h_{j})/\sigma$; these visualize tail-risk (rare catastrophic flips) that raw accuracy masks and make robustness audits reproducible.
7  Related Work
There have been multiple benchmarks for evaluating the mathematical-reasoning capabilities of large language models (LLMs). Early math‐reasoning benchmarks such as MATH(1.25 k problems) (Hendrycks et al., 2021), and GSM8K(8.5 k problems) (Cobbe et al., 2021), revealed basic arithmetic/algebra skills. But their difficulty is now saturated as LLMs scale. For instance, with prompting strategies such as DUP, GPT-4 attains 97.1% accuracy on GSM8K (Zhong et al., 2025). This ceiling at the high-school-competition level motivated the creation of a new generation of harder benchmarks.
Subsequent benchmarks target harder problems. Omni-MATH contributes 4 428 rigorously annotated Olympiad-level problems (Gao et al., 2024). Likewise, OlympiadBench provides a bilingual, multimodal benchmark of 8 476 Olympiad-level math and physics problems with expert step-by-step solutions (He et al., 2024). The cross-disciplinary benchmark ARB consist questions in mathematics, physics, biology, chemistry, and law, with a rubric-based self-grading protocol (Sawada et al., 2023). Some other benchmarks focuses specifically on formal proof. MiniF2F supplies 488 Olympiad-level problems formalized in multiple proof assistants (Zheng et al., 2022). Putnambench, offers 1 692 rigorously hand-crafted formalizations of Putnam Competition problems (Tsoukalas et al., 2024).
Nevertheless, recent studies warn that scores on many NLP benchmarks may be artificially inflated by data contamination, when LLMs are trained on the benchmark questions. Sainz et al. (2023) point out that many benchmarks may be inflated because large language models often memorize test data seen during pre-training.  Balloccu et al. (2024) conduct a systematic audit of data leakage for closed-source LLMs and estimate that roughly 4.7 million test examples from 263 datasets were likely exposed to the models.
In order to obtain a more robust evaluation of LLMs’ reasoning capabilities, it is important to prevent data leakage. One approach is to create original questions. FrontierMath, for example, addresses this with a rigorously curated benchmark of hundreds of original, expert-level mathematics problems spanning fields from number theory to algebraic geometry (Glazer et al., 2024). Putnam-AXIOM takes this approach and comprises 522 challenging problems from the William Lowell Putnam Competition plus 100 programmatically generated functional variations, furnishing a contamination-resilient benchmark (Gulati et al., 2025).
Another method to deal with data leakage is introducing contrast sets-small, label-changing perturbations of existing test instances—to probe a model’s local decision boundary (Gardner et al., 2020). Huang et al. (2025) construct MATH-Perturb that applies simple and hard perturbations to 279 level-5 MATH problems and discover that models suffer 12–16 percentage-point drops on the hard variant. Shalyt et al. (2025) complement this line of work with ASyMOB, a 17k-problem benchmark whose algebra-focused numerical and symbolic perturbations reveal performance drops of up to 70 percentage points, highlighting models’ fragility under such stress tests. Similarly, Yu et al. (2025) propose Math-RoB, a synthetic benchmark that enables robustness evaluation against the data contamination through an instruction-based approach. These efforts either focus on specific aspects that limit generalizability, are based on benchmarks that are too easy for current models, or introduce transformations that are not mathematically equivalent, thus confounding true robustness evaluation.
Building on these prior efforts, our work introduces Generalization–and–Perturbation (GAP), a unified framework that addresses both data leakage and robustness by generating mathematically equivalent variants of complex problems, significantly expanding the evaluation depth of existing benchmarks. The framework can be applied to existing and future benchmarks, and all types of questions to strengthen their reliability. To address the saturation of accuracy scores, we apply our framework to challenging, college-level competition mathematics problems. We instantiate GAP on every William Lowell Putnam Competition problem from 1938–2024 (1 051 originals), expanding each item into five mathematically equivalent variants and thereby producing PutnamGAP, a corpus of 6,306 stress-test questions. Finally, we release an open-source evaluation stack that rigorously grades solutions step by step, making assessment fully automated, transparent, and reproducible.
8  Conclusion & Future Work
Robust reasoning is required in many applications of LLMs. In this paper, we proposed a novel Generalization–and–Perturbation (GAP) framework for analyzing and evaluating robustness of LLMs’ reasoning capacity. By instantiating GAP on all 1,051 Putnam Competition questions we produced the 6,306‑question PutnamGAP benchmark. A zero‑shot evaluation of 18 commercial and open‑source LLMs revealed sharp and consistent accuracy drops. These results expose a clear robustness gap that leaderboard scores on unperturbed datasets have so far not shown.
Our findings highlight three actionable directions.
  • Benchmarking: GAP offers an open‑ended supply of contamination‑resistant test items, limiting future data leakage and score inflation.
  • Training: curricula that randomize both symbol identities and numeric parameters during fine‑tuning should become standard practice for models targeting formal reasoning domains.
  • Security: the same surface‑level fragility that hurts accuracy can be weaponized for prompt‑injection attacks, so GAP‑style mutation should be built into red‑teaming pipelines.
There are multiple interesting future research directions based on our work: (i) diversify the verifier ensemble with symbolic provers and heterogeneous LLMs to rule out collusive blind spots, (ii) port GAP to applied mathematics, physics and multi-modal STEM corpora, and (iii) integrate on‑the‑fly GAP transformations into training so that invariance to symbol and parameter changes is learned rather than merely tested.
PutnamGAP makes one lesson unmistakable: genuine progress in mathematical AI will be measured not by ever‑higher raw scores, but by a model’s ability to stride across the hidden gulf between symbols and substance. The next generation of top‑tier systems will earn their place only by refusing to be left behind on GAPs.